# Hierarchical filtrations of torsion-free sheaves and birational geometry

### v1

@@ Abstract.
We introduce the notion of hierarchical filtrations of torsion-free sheaves on normal projective varieties and define the associated numerical invariant called hierarchical depth. This invariant measures the maximal length of filtrations by saturated subsheaves of equal rank whose successive quotients are torsion sheaves supported in codimension one.
We establish general bounds for hierarchical depth in terms of the divisor class of the determinant and give exact formulas in several basic geometric situations, including the case of smooth projective curves and varieties of Picard rank one. A key technical ingredient is the study of elementary transforms along effective divisors and their commutativity properties.
In dimension two, we analyze the behavior of hierarchical depth under birational morphisms and show that it admits a precise description along the minimal model program. In particular, we prove that hierarchical depth transforms additively with respect to exceptional divisors and is explicitly computable on minimal models.
As an application, we relate hierarchical depth to degeneracies in algebraic–geometric codes and show that birational simplification via the minimal model program leads to effective improvements of code parameters. These results demonstrate that hierarchical depth provides a new bridge between the birational geometry of vector bundles and arithmetic applications.
@@ Introduction
Filtrations of vector bundles and coherent sheaves are among the most classical tools in algebraic geometry. Since the foundational work of Grothendieck on coherent sheaves and their cohomology [4], filtrations have provided a systematic way to decompose global objects into simpler pieces reflecting both geometric and numerical properties. Standard references such as [5, 3, 20, 21] illustrate the central role of torsion-free and reflexive sheaves in the study of divisors, intersection theory, and birational geometry.
One of the most influential filtration-theoretic constructions is the Harder–Narasimhan filtration, originally introduced for vector bundles on curves [7] and later extended to higher-dimensional varieties. This filtration underlies much of modern stability theory and plays a fundamental role in the geometry of moduli spaces of sheaves, as developed for instance in [12, 13, 8]. Closely related are Jordan–Hölder filtrations and refinements appearing in the theory of semistable sheaves. These constructions, however, depend essentially on the choice of a polarization and are therefore sensitive to auxiliary data.
In parallel, birational geometry has undergone a profound development through the minimal model program. Initiated by Mori and Shokurov [17, 18, 19], and culminating in the existence of minimal models for varieties of log general type [1], the MMP provides a structural framework for understanding higher-dimensional algebraic varieties via divisorial contractions, flips, and minimal models. Modern expositions may be found in [2, 10]. Despite this progress, comparatively little is known about how intrinsic invariants of vector bundles and torsion-free sheaves behave along the steps of the minimal model program.
@@
The purpose of this paper is to introduce a new filtration-theoretic framework for torsion-free sheaves that is intrinsically adapted to divisorial and birational geometry. We define the notion of a hierarchical filtration, consisting of successive saturated subsheaves of equal rank whose quotients are torsion sheaves supported in codimension one. Associated to such filtrations is a discrete numerical invariant, called the hierarchical depth, measuring the maximal possible length of such a filtration. Hierarchical filtrations were first introduced and studied for line bundles in [14], where they were applied to the construction of optimal algebraic–geometric codes. The present paper develops a substantially broader theory for torsion-free sheaves of arbitrary rank and investigates its interaction with birational geometry and the minimal model program. Unlike stability-theoretic invariants, hierarchical depth is independent of any choice of polarization and is governed instead by the determinant of the sheaf and the divisor theory of the underlying variety.
The first part of the paper develops the basic properties of hierarchical filtrations. Using standard techniques from the theory of divisors and intersection theory [3, 9], we prove finiteness of hierarchical depth and derive effective bounds in terms of the divisor class of the determinant. In several basic geometric situations, including smooth projective curves and varieties of Picard rank one, we obtain explicit formulas. These results show that hierarchical depth is a computable invariant determined by classical numerical data.
A key technical aspect of the theory is the interpretation of hierarchical filtrations via iterated elementary transforms along effective divisors. Elementary transforms appear naturally in the study of vector bundles and reflexive sheaves, particularly in connection with degenerations and modifications along divisors; see for instance [6, 13]. We show that, under suitable conditions, such transforms commute and may be reordered without affecting the resulting hierarchical depth. This yields a concrete description of the invariant and clarifies its independence from auxiliary choices.
The second part of the paper is devoted to the interaction between hierarchical depth and birational geometry. After establishing general functorial properties under pullback and pushforward along birational morphisms, we specialize to the case of surfaces. In dimension two, we prove that hierarchical depth transforms additively along the steps of the minimal model program. In particular, the invariant admits an explicit description on minimal models, with its variation controlled by exceptional divisors and their numerical contributions. This situates hierarchical depth as a genuinely birational invariant in the surface case, in the spirit of the MMP as developed in [1, 2, 10].
Finally, we present applications to algebraic–geometric codes constructed from vector bundles. Vector bundle methods in coding theory have been studied by several authors; see for example [11, 16, 15]. Building on earlier work on hierarchical filtrations for line bundles [14], we show that hierarchical depth detects divisorial degeneracies affecting code parameters. By combining the filtration theory developed here with birational simplification via the minimal model program, we obtain effective improvements in code constructions.
The paper is organized as follows. Section 1 reviews preliminaries and fixes notation. Section 2 introduces hierarchical filtrations and hierarchical depth. Section 3 is devoted to elementary transforms and their commutativity properties. Determinantal bounds and finiteness results are established in Section 4. In Section 5 we analyze the behavior of hierarchical depth under birational morphisms and study its interaction with the surface minimal model program. Slope behavior under increments, birational stability under the MMP and comparisons with classical filtrations are given in Section 6. Applications to algebraic–geometric codes are discussed in Section 7.
@@ Acknowledgments
1. Preliminaries and conventions
Throughout this paper, we work over an algebraically closed field $k$ of arbitrary characteristic. All schemes are assumed to be separated and of finite type over $k$.
1.1. Varieties and divisors
A variety is an integral normal projective scheme over $k$. Unless explicitly stated otherwise, all varieties considered in this paper have dimension at most $2$.
For a variety $X$, we denote by $\mathrm{Pic}(X)$ its Picard group and by $\mathrm{Div}(X)$ the group of Weil divisors on $X$. A divisor $D\in\mathrm{Div}(X)$ is called effective, denoted $D\geq 0$, if it is a finite non-negative linear combination of irreducible prime divisors. Linear equivalence of divisors is denoted by $\sim$.
When $X$ is smooth, Weil and Cartier divisors are identified. In general, divisor classes will be considered in the divisor class group of $X$.
1.2. Coherent sheaves and saturation
All sheaves in this paper are coherent $\mathcal{O}_{X}$-modules. A coherent sheaf $\mathcal{E}$ on $X$ is called torsion-free if it has no nonzero subsheaves supported in codimension at least $1$.
Let $\mathcal{E}$ be a torsion-free sheaf on $X$. A subsheaf $\mathcal{F}\subseteq\mathcal{E}$ is said to be saturated if the quotient $\mathcal{E}/\mathcal{F}$ is torsion-free. Equivalently, $\mathcal{F}$ is saturated if it coincides with its image in the reflexive hull $\mathcal{E}^{\vee\vee}$.
Standing convention. Unless explicitly stated otherwise, all subsheaves of torsion-free sheaves are assumed to be saturated.
With this convention, if $\mathcal{F}\subseteq\mathcal{E}$ are torsion-free sheaves of the same rank, then the quotient $\mathcal{E}/\mathcal{F}$ is a torsion sheaf supported in codimension $1$.
1.3. Determinants
Let $\mathcal{E}$ be a torsion-free sheaf of rank $r$ on a normal variety $X$. The determinant of $\mathcal{E}$ is defined by
@@ 1.3. Determinants
which is a rank-one reflexive sheaf. We identify $\det(\mathcal{E})$ with the corresponding divisor class on $X$.
If
$0\longrightarrow\mathcal{F}\longrightarrow\mathcal{E}\longrightarrow\mathcal{Q}\longrightarrow 0$
is an exact sequence of coherent sheaves on $X$ with $\mathcal{F}$ and $\mathcal{E}$ torsion-free of the same rank, then $\mathcal{Q}$ is a torsion sheaf supported in codimension $1$, and there exists an effective divisor $D$ on $X$ such that
@@
We refer to $D$ as the determinantal divisor associated with the inclusion $\mathcal{F}\subseteq\mathcal{E}$.
1.4. Elementary transforms
Let $\mathcal{E}$ be a torsion-free sheaf on $X$, and let $D$ be an effective divisor on $X$. An elementary transform of $\mathcal{E}$ along $D$ is a torsion-free sheaf $\mathcal{E}^{\prime}$ fitting into an exact sequence
@@ 1.4. Elementary transforms
where $\mathcal{T}$ is a torsion sheaf supported on $D$.
Elementary transforms preserve the rank of the sheaf and modify the determinant by an effective divisor supported on $D$. They will be used extensively in the construction of hierarchical filtrations.
@@ 2. Hierarchical filtration of vector bundles
Let $X$ be a smooth projective variety over a field. Fix a reference line bundle
$\Lambda_{0}\in\mathrm{Pic}(X)$
(which most commonly one takes to be $\mathcal{O}_{X}$). Let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$.
A hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$ is a finite chain of inclusions of locally free $\mathcal{O}_{X}$-subsheaves
@@
satisfying the following conditions for every $i=1,\dots,h$:
  1. (1)
    each $\mathcal{E}_{i}$ is a vector bundle of rank $r$, and the inclusion $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is saturated, i.e. the quotient
    $\mathcal{Q}_{i}:=\mathcal{E}_{i}/\mathcal{E}_{i-1}$
    is a nonzero torsion sheaf (equivalently $\mathcal{E}_{i-1}$ is a saturated subsheaf of $\mathcal{E}_{i}$).
  2. (2)
    there exists a nonzero effective Cartier divisor $D_{i}\subset X$ such that the following equivalent formulation holds:
  3. (3)
    the filtration is normalized by the requirement
@@
Note that each one of two following statements are equivalent to condition (2) of above definition:
(2-1) the torsion sheaf $\mathcal{Q}_{i}$ is supported (set-theoretically) on $D_{i}$, and
$c_{1}(\mathcal{E}_{i})=c_{1}(\mathcal{E}_{i-1})+[D_{i}]\quad\in\mathrm{NS}(X).$
(2-2) locally on $X$ the inclusion $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is obtained from $\mathcal{E}_{i-1}$ by an elementary transform (Hecke modification) along the Cartier divisor $D_{i}$; equivalently there is an exact sequence
@@
    Equivalence of (2),(2-1),(2-2). Under the saturation hypothesis the three formulations in (2)(a)–(c) are equivalent: (c) implies (b) by taking first Chern classes and (b) implies (a) because the determinant of a short exact sequence with torsion quotient acquires the line bundle associated to the divisor supporting the torsion; conversely (a) together with saturation forces the quotient to be torsion supported on the divisor appearing in the determinant relation.
@@ 4. Hierarchical depth and determinant bounds
Proposition 4.3.
Let $X$ be a smooth projective variety, fix a reference line bundle $\Lambda_{0}\in\mathrm{Pic}(X)$, and let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$.
@@
    Suppose $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a saturated inclusion of locally free sheaves of the same rank $r$, so that $\mathcal{E}/\mathcal{E}^{\prime}$ is a nonzero torsion sheaf. Assume $\det(\mathcal{E}^{\prime})$ is measured relative to the same normalization $\Lambda_{0}$ as $\det(\mathcal{E})$. Then
@@
    If $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a subbundle of strictly smaller rank $\mathrm{rank}(\mathcal{E}^{\prime})<\mathrm{rank}(\mathcal{E})$, no inequality between $h_{\Lambda_{0}}(\mathcal{E}^{\prime})$ and $h_{\Lambda_{0}}(\mathcal{E})$ holds in general. In particular, one may have $h_{\Lambda_{0}}(\mathcal{E}^{\prime})>h_{\Lambda_{0}}(\mathcal{E})$.
@@
    Let
    $0\longrightarrow\mathcal{E}^{\prime}\longrightarrow\mathcal{E}\longrightarrow\mathcal{E}^{\prime\prime}\longrightarrow 0$
    be a short exact sequence of vector bundles. Assume that:
    1. (i)
      the hierarchical filtration of $\mathcal{E}^{\prime}$ is normalized by $\Lambda_{0}$;
    2. (ii)
      the hierarchical filtration of $\mathcal{E}^{\prime\prime}$ is normalized by $\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})$; and
    3. (iii)
      every determinant increment divisor of the $\mathcal{E}^{\prime\prime}$–filtration lifts to a saturated same–rank elementary transform of $\mathcal{E}$ (equivalently: the effective divisors involved in the filtration of $\mathcal{E}^{\prime\prime}$ are disjoint from the support of the torsion in successive pushforwards during the reconstruction of $\mathcal{E}$).
    Under these hypotheses the concatenation of hierarchical filtrations yields a hierarchical filtration of $\mathcal{E}$, and therefore
    $h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})}(\mathcal{E}^{\prime\prime}),$
    and in particular
    (3) $h_{\Lambda_{0}}(\mathcal{E})\geq\max\Big\{h_{\Lambda_{0}}(\mathcal{E}^{\prime}),h_{\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})}(\mathcal{E}^{\prime\prime})\Big\}.$
  4. (D)
    If $\mathcal{E}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}$, and the determinant increments appearing in maximal hierarchical filtrations of $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ admit representatives whose divisors have disjoint support (so that the corresponding elementary transforms commute), then their interleaving produces a hierarchical filtration of $\mathcal{E}$ and
    $h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}_{1})+h_{\Lambda_{0}}(\mathcal{E}_{2}).$
    Without the disjoint–support (or commutativity) hypothesis the inequality may be strict or may fail in the naive additive form.
@@
(A) A hierarchical filtration of $\mathcal{E}^{\prime}$ consists of a sequence of saturated same–rank inclusions
$\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{h^{\prime}}=\mathcal{E}^{\prime},$

### v3

# Hierarchical filtrations of vector bundles and birational geometry

@@ Abstract.
We introduce and systematically study hierarchical filtrations of vector bundles on smooth projective varieties. These are filtrations by saturated subsheaves of equal rank whose successive quotients are torsion sheaves supported in codimension one. The associated numerical invariant, called hierarchical depth, measures the maximal length of such filtrations.
We establish general bounds for hierarchical depth in terms of the determinant class and provide exact formulas for smooth curves and varieties of Picard rank one. A key technical result concerns the commutativity of elementary transforms along disjoint divisors and their role in constructing filtrations.
For surfaces, we analyze the behavior of hierarchical depth under birational morphisms and prove that it transforms additively along the steps of the minimal model program. In particular, we obtain an explicit formula relating the depth on a surface to that on its minimal model via exceptional divisor contributions.
As an application, we connect hierarchical depth to degeneracies in algebraic–geometric codes and show that birational simplification via the MMP leads to effective improvements of code parameters. This establishes hierarchical depth as a new bridge between birational geometry, vector bundle theory, and coding theory.
@@ Introduction
Filtrations of vector bundles are fundamental tools in algebraic geometry, from Grothendieck’s foundational work on coherent sheaves [4] to the Harder–Narasimhan filtration in stability theory [7]. While stability filtrations depend on a choice of polarization, there remains a need for intrinsic, polarization–free filtrations that reflect divisorial and birational properties.
@@
This paper introduces a new filtration–theoretic framework for torsion–free sheaves adapted precisely to birational geometry. We define hierarchical filtrations—chains of saturated subsheaves of equal rank with torsion quotients supported in codimension one—and their associated numerical invariant, the hierarchical depth. Unlike stability–theoretic invariants, hierarchical depth is independent of any polarization and governed instead by the determinant and the divisor theory of the underlying variety.
Our work builds on and substantially extends the hierarchical filtration theory for line bundles developed in [14], where applications to algebraic–geometric codes were first explored. The present paper develops a comprehensive theory for arbitrary rank, investigates its interaction with birational geometry and the minimal model program, and establishes new connections to coding theory.
Main results
  1. 1.
    We introduce hierarchical filtrations and hierarchical depth (Definition 2.1), prove their finiteness (Proposition 2.4), and give explicit formulas in basic geometric situations (Propositions 4.1, 2.4).
  2. 2.
    We establish commutativity properties for elementary transforms along disjoint divisors (Lemma 3.2, Corollary 3.4), providing a concrete mechanism for constructing filtrations.
  3. 3.
    For surfaces, we prove that hierarchical depth transforms additively under birational morphisms arising in the MMP (Theorem 5.6). In particular, it admits an explicit description on minimal models with variation controlled by exceptional divisors.
  4. 4.
    We relate hierarchical depth to slope stability (Theorem 6.1) and show it is preserved under birational morphisms (Corollary 6.3).
  5. 5.
    As applications to coding theory, we demonstrate that hierarchical depth detects divisorial degeneracies affecting algebraic–geometric code parameters (Proposition 7.1). Birational simplification via the MMP leads to effective improvements in both absolute and normalized minimum distances (Corollary 7.5)
These results situate hierarchical depth as a genuinely birational invariant in the surface case and provide a new computational tool linking the geometry of vector bundles to arithmetic applications.
@@ Acknowledgments
1. Preliminaries
Conventions. Throughout, $\mathbb{F}$ denotes an algebraically closed field of arbitrary characteristic. All schemes are separated and of finite type over $\mathbb{F}$. A variety is an integral normal projective $\mathbb{F}$-scheme. Unless otherwise stated, varieties are assumed smooth; the restriction to dimension at most $2$ applies only when invoking the minimal model program in Sections 5 and onward.
Divisors and linear equivalence. For a variety $X$, denote by $\mathrm{Div}(X)$ the group of Weil divisors and by $\mathrm{Pic}(X)$ the Picard group. When $X$ is smooth, Cartier and Weil divisors coincide. An effective divisor is written $D\geq 0$. Linear equivalence is denoted by $\sim$.
Sheaves and determinants. All sheaves are coherent $\mathcal{O}_{X}$-modules. A coherent sheaf $\mathcal{E}$ is torsion-free if it has no nonzero subsheaves supported in codimension $\geq 1$. A subsheaf $\mathcal{F}\subseteq\mathcal{E}$ is saturated if $\mathcal{E}/\mathcal{F}$ is torsion-free.
For a torsion-free sheaf $\mathcal{E}$ of rank $r$, its determinant is
@@ 1. Preliminaries
identified with the corresponding divisor class. If $\mathcal{F}\hookrightarrow\mathcal{E}$ is an inclusion of torsion-free sheaves of equal rank with torsion quotient, there exists an effective divisor $D$ (the determinantal divisor) such that
@@
Elementary transforms. Let $\mathcal{E}$ be a torsion-free sheaf on $X$ and $D\subset X$ an effective divisor. An elementary transform of $\mathcal{E}$ along $D$ is a sheaf $\mathcal{E}^{\prime}$ fitting into
@@
where $\mathcal{T}$ is a torsion sheaf supported on $D$. Such transforms preserve rank and satisfy $\det(\mathcal{E}^{\prime})\simeq\det(\mathcal{E})(-D)$.
Birational morphisms. For a proper birational morphism $f:Y\to X$ between normal varieties:
  • The pullback $f^{*}\mathcal{E}$ of a torsion-free sheaf $\mathcal{E}$ on $X$ may contain torsion; in practice we work with its torsion-free quotient.
  • The pushforward $f_{*}\mathcal{F}$ of a torsion-free sheaf $\mathcal{F}$ on $Y$ is understood to be replaced by its saturation when viewed as a subsheaf on $X$.
These conventions ensure that all sheaves considered remain torsion-free and all inclusions saturated.
Cohomology and slopes. We write $h^{i}(X,\mathcal{F})=\dim_{\mathbb{F}}H^{i}(X,\mathcal{F})$. For a smooth projective variety $X$ with ample divisor $H$, the slope of a torsion-free sheaf $\mathcal{E}$ is
$\mu_{H}(\mathcal{E})=\frac{c_{1}(\det\mathcal{E})\cdot H^{\dim X-1}}{\mathrm{rank}\mathcal{E}}.$
@@ 2. Hierarchical filtration of vector bundles
Let $X$ be a smooth projective variety over a field, $\Lambda_{0}\in\mathrm{Pic}(X)$ a reference line bundle (typically $\Lambda_{0}=\mathcal{O}_{X}$), and $\mathcal{E}$ a vector bundle of rank $r$ on $X$. A hierarchical filtration of $\mathcal{E}$ normalized by $\Lambda_{0}$ is a finite chain
@@
where for each $i=1,\dots,h$:
  1. (i)
    $\mathcal{E}_{i}$ is a vector bundle of rank $r$, and $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is saturated;
  2. (ii)
    There exists a nonzero effective Cartier divisor $D_{i}$ with $\det(\mathcal{E}_{i})\simeq\det(\mathcal{E}_{i-1})\otimes\mathcal{O}_{X}(D_{i})$;
  3. (iii)
@@
Note that each one of two following statements are equivalent to condition (ii) of above definition:
(ii-1) the torsion sheaf $\mathcal{Q}_{i}$ is supported (set-theoretically) on $D_{i}$, and
$c_{1}(\mathcal{E}_{i})=c_{1}(\mathcal{E}_{i-1})+[D_{i}]\ \in\mathrm{NS}(X).$
(ii-2) locally on $X$ the inclusion $\mathcal{E}_{i-1}\hookrightarrow\mathcal{E}_{i}$ is obtained from $\mathcal{E}_{i-1}$ by an elementary transform (Hecke modification) along the Cartier divisor $D_{i}$; equivalently there is an exact sequence
@@ 4. Hierarchical depth and determinant bounds
Proposition 4.3 (Basic properties of hierarchical depth).
Let $X$ be a smooth projective variety, $\Lambda_{0}\in\mathrm{Pic}(X)$ a reference line bundle, and $\mathcal{E}$ a vector bundle of rank $r$ on $X$.
@@
    (Monotonicity) If $\mathcal{E}^{\prime}\subset\mathcal{E}$ is a saturated inclusion of vector bundles of the same rank $r$, then
@@
    (Extension inequality) Let $0\to\mathcal{E}^{\prime}\to\mathcal{E}\to\mathcal{E}^{\prime\prime}\to 0$ be a short exact sequence of vector bundles on $X$.
    Suppose there exist hierarchical filtrations
    $\mathfrak{F}^{\prime}:\mathcal{E}^{\prime}_{0}\subset\mathcal{E}^{\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime}_{a}=\mathcal{E}^{\prime},\qquad\mathfrak{F}^{\prime\prime}:\mathcal{E}^{\prime\prime}_{0}\subset\mathcal{E}^{\prime\prime}_{1}\subset\cdots\subset\mathcal{E}^{\prime\prime}_{b}=\mathcal{E}^{\prime\prime},$
    normalized by $\Lambda_{0}$ and $\Lambda_{0}^{\prime}:=\Lambda_{0}\otimes\det(\mathcal{E}^{\prime})^{-1}$ respectively. Assume that for each $j=1,\dots,b$, the elementary transform $\mathcal{E}^{\prime\prime}_{j-1}\subset\mathcal{E}^{\prime\prime}_{j}$ along an effective divisor $D_{j}$ lifts to an elementary transform of $\mathcal{E}$ in the following sense: there exists a commutative diagram with exact rows
    ${0}{\mathcal{E}^{\prime}}$$\mathcal{F}_{j-1}$$\mathcal{E}^{\prime\prime}_{j-1}$$0{0}{\mathcal{E}^{\prime}}$$\mathcal{F}_{j}$$\mathcal{E}^{\prime\prime}_{j}$}0{0}
    where:
    • $\mathcal{F}_{j-1},\mathcal{F}_{j}\subset\mathcal{E}$ are saturated subbundles of rank $\mathrm{rank}(\mathcal{E})$,
    • the vertical inclusions $\mathcal{F}_{j-1}\hookrightarrow\mathcal{F}_{j}$ and $\mathcal{E}^{\prime\prime}_{j-1}\hookrightarrow\mathcal{E}^{\prime\prime}_{j}$ are elementary transforms along $D_{j}$,
    • the middle vertical map induces the given inclusion $\mathcal{E}^{\prime\prime}_{j-1}\hookrightarrow\mathcal{E}^{\prime\prime}_{j}$ on the quotients.
    Then
    $h_{\Lambda_{0}}(\mathcal{E})\geq h_{\Lambda_{0}}(\mathcal{E}^{\prime})+h_{\Lambda^{\prime}_{0}}(\mathcal{E}^{\prime\prime}).$
@@
    (Direct sums) If $\mathcal{E}=\mathcal{E}_{1}\oplus\mathcal{E}_{2}$ and there exist maximal hierarchical filtrations of $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ whose determinant increments have pairwise disjoint support, then
    $h_{\Lambda_{0}}(\mathcal{E})=h_{\Lambda_{0}}(\mathcal{E}_{1})+h_{\Lambda_{0}}(\mathcal{E}_{2}).$
    Without disjoint supports, only the inequality $\geq$ holds in general.